cos(x-y) = -1 show that cosx+ cos y=0 and sinx + siny=0

cos(x-y) = -1  show that cosx+ cos y=0  and sinx + siny=0

Grade:12

1 Answers

Prudhvi teja
13 years ago

Dear Manoj

cos(x-y) = -1 

that means

x-y = (2n+1)pi

x=y+(2n+1)pi

apply cos on both sides

cos x = cos(y+(2n+1)pi)

=> cosx = -cosy

=> cos x + cos y = 0

in the same way sinx+siny = 0

 

Please feel free to post as many doubts on our discussion forum as you can.
we will get you the answer and detailed  solution very  quickly.

All the best.
 
Regards,
Askiitians Experts
Prudhvi Teja